Calculate the percent ionization of hypochlorous acid (HClO) in solutions of each of the following concentrations (Ka = 3.0e-08.)



(a) 0.144 M

%



(b) 0.568 M

%



(c) 0.896 M

%
Answer #1

a)
HClO dissociates as:

HClO          ----->     H+   + ClO-
0.144                 0         0
0.144-x               x         x


Ka = [H+][ClO-]/[HClO]
Ka = x*x/(c-x)
Assuming x can be ignored as compared to c
So, above expression becomes

Ka = x*x/(c)
so, x = sqrt (Ka*c)
x = sqrt ((3*10^-8)*0.144) = 6.573*10^-5

since c is much greater than x, our assumption is correct
so, x = 6.573*10^-5 M



% dissociation = (x*100)/c
= 6.573*10^-5*100/0.144
= 4.564*10^-2 %
Answer: 4.6*10^-2 %

b)
HClO dissociates as:

HClO          ----->     H+   + ClO-
0.568                 0         0
0.568-x               x         x


Ka = [H+][ClO-]/[HClO]
Ka = x*x/(c-x)
Assuming x can be ignored as compared to c
So, above expression becomes

Ka = x*x/(c)
so, x = sqrt (Ka*c)
x = sqrt ((3*10^-8)*0.568) = 1.305*10^-4

since c is much greater than x, our assumption is correct
so, x = 1.305*10^-4 M



% dissociation = (x*100)/c
= 1.305*10^-4*100/0.568
= 2.298*10^-2 %
Answer: 2.3*10^-2 %

c)

HClO dissociates as:

HClO          ----->     H+   + ClO-
0.896                 0         0
0.896-x               x         x


Ka = [H+][ClO-]/[HClO]
Ka = x*x/(c-x)
Assuming x can be ignored as compared to c
So, above expression becomes

Ka = x*x/(c)
so, x = sqrt (Ka*c)
x = sqrt ((3*10^-8)*0.896) = 1.64*10^-4

since c is much greater than x, our assumption is correct
so, x = 1.64*10^-4 M



% dissociation = (x*100)/c
= 1.64*10^-4*100/0.896
= 1.83*10^-2 %
Answer: 1.8*10^-2 %